# Sales Whisperer: A Human-Inconspicuous Attack on LLM Brand Recommendations

### v1

@@ 6.1 Observations on Paraphrased Prompts
As we described in §5.2, each of the product categories in our dataset has three to ten paraphrased prompts. For each prompt, we measured the probability that LLMs would mention some target brand (described in §5.2) of a category within the first 64 tokens of the response. For each combination of product category and brand, we computed the maximal difference in the probability of responses mentioning the target brand between paraphrased prompts of the product category (within the first tokens). Adversaries may paraphrase a low-probability base prompt in order to find the high-probability prompts to achieve the goals described in §3.
@@
We summarize our results in Fig. 6. Specifically, on four open-sourced LLMs—Gemma-it (instruction-tuned), Llama2, Llama3, and Llama3-it (instruction-tuned)—the likelihood of the target brand being mentioned in the responses can differ between a pair of paraphrased prompts by up to , , with the average of this absolute difference being , , , and , respectively. For example, when comparing responses to “I’m curious to know your preference for the pressure cooker that offers the best combination of cooking performance, durable construction, and overall convenience in preparing a variety of dishes.” with the prompt paraphrased as “Can you recommend the ultimate pressure cooker that excels in providing consistent pressure, user-friendly controls, and additional features such as multiple cooking presets or a digital display for precise settings?”, the probability of Gemma-it mentioning the brand “InstantPot” (“pressure cooker” product category) within the first tokens of the response went from to .
@@
Our results suggest that while paraphrased prompts appear similar to humans the responses to the prompts can differ substantially in how likely they are to mention a target brand (within the first tokens of the response). Therefore, an adversary wanting to promote a certain brand can use this to their advantage: The adversary may try various paraphrases of prompts, test the prompts, and pick the paraphrase that results in the highest probability of the target brand being mentioned, ultimately promoting the brand when prompt is used in the real world. While we generated these paraphrases using ChatGPT, and confirmed that they were valid and reasonable, adversaries may also be able to create paraphrases manually or by another method, and may be able to test even more paraphrases than we did in our measurements.
6.2 Synonym-Replaced Adversarial Prompts
@@
As described in §5.2, we measured and now report how often our synonym-replacement approach improves the probability of a brand being mentioned. As opposed to paraphrasing, in the synonym-replacement approach we automatically generate a set of potential candidate prompts by perturbing a base prompt via synonym replacement, without needing to confirm whether these perturbed prompts are valid. The prompt with the lowest loss is selected, and we assess how well these selected prompts increase the probability that the target brand is mentioned. We evaluate the average improvement over multiple models, at different response lengths, and for different probabilities of a brand being mentioned in the base prompt. In addition, other candidate prompts could also be tested, although we did not take this approach in this paper. We used loss as a metric that narrowed down the large set of potential perturbed prompts we found using synonym replacement to one. Therefore, we are interested in exploring the highest improvement we can achieve between a base and perturbed score, as adversaries would be able to use a similar method to §6.1 and explore multiple perturbated prompts. So, we also describe the largest increases in probability of mentioning a target brand in the responses to the perturbed prompt we found via our synonym-replacement method compared to the base prompt. Our evaluation shows that our synonym-replacement attack increases the likelihood of LLMs mentioning a brand on average.
For all evaluations, we focus on the first 64 tokens of the response (see subsection 5.1 for details).
@@
The results for average improvement in likelihoods of a brand being mentioned in the responses to perturbed prompts created via synonym replacement compared to in response to the original base prompt, across models, are shown in Fig. 7. Overall, we find that our attack results in an average absolute improvement in all models. Specifically, for brands that were mentioned at least once (i.e., of the time) in responses to the base prompt, Gemma-it had an average absolute improvement within tokens and a improvement within tokens. Llama2 achieves an average absolute improvement of within tokens, and within tokens. Llama3 achieves an average absolute improvement of within tokens, and within tokens. Llama3-it achieves an average absolute improvement of within tokens, and within tokens. The averages of the best absolute improvement in each product category are , , , and on Gemma-it, Llama2, Llama3, and Llama3-it respectively. Our synonym replacement approach achieves a positive absolute improvement in probabilities, which verifies that our approach is capable of forcing LLMs to mention target brands more often within some number of tokens. On all four models, we see a bigger absolute improvement within some number of tokens less than . While we only collected responses of length 64 tokens for our 1,809 attacks for each model (see subsection 5.1), we did run a select set of base and perturbed prompt pairs (chosen in §5.2) to full completion. While we did not see improvements in all pairs (for example, some base prompts had near probabilities of mentioning certain brands with long completions, not allowing for any improvement), we still saw pairs with significant increases when considering long responses. For Llama3-it, the perturbed prompt “Can you recommend the superior video game console that excels in providing top-notch graphics, dissimilar gaming options, and additional features such as online connectivity, appropriate for both casual and hardcore gamers?” was more likely than its base prompt “Can you recommend the ultimate video game console that excels in providing top-notch graphics, diverse gaming options, and additional features such as online connectivity, suitable for both casual and hardcore gamers?” (i.e., “ultimate” was changed to “superior” and “diverse” was changed to “dissimilar”) to mention Xbox in long completions, even more than when completions were only 64 tokens long (). We evaluate if attack objectives are met with full responses in a more realistic setting in §7.2.2.
Baseline probability and attack success
Here we explore the relationship between attack success rates and the probability of a brand being mentioned in responses to the base prompt. As shown in Fig. 8, our synonym-replacement approach on average achieved a bigger absolute improvement when the target brand was mentioned more often in responses to the base prompt before perturbation. When the target brands were mentioned at least (i.e., in out of responses before perturbation), our synonym-replacement approach achieved an absolute improvement of , and on Gemma-it, Llama2, Llama3 and Llama3-it, respectively. We see that for base prompts with higher probabilities of mentioning the target brand, our perturbed prompts, on average, have a higher increase in probability. This indicates that our perturbation method is more effective on an identifiable domain, i.e. prompts where the target brand is mentioned more or less frequently, than it is on the overall average case.
@@
Besides the average absolute improvement, we also explored the maximum absolute improvement among all combinations of base prompts and brands, as shown in Fig. 9. While these results don’t represent the expected improvement using this method, they do demonstrate that it is possible to find pairs of prompts with vastly different probabilities of mentioning a certain brand. In §6.1 we showed this for prompts that were manually paraphrased; here, we show that the synonym-replacement attack can find such alternative prompts automatically. Further, the prompts generated by synonym-replacement differ from their base prompts minimally (at most a seven synonym difference in our dataset), and are perceptually the same along multiple dimensions (see §7.2.2).
We were able to achieve bigger absolute improvement on Gemma-it compared to the other three Llama models, although Gemma-it does not always have the highest average absolute improvement (Fig. 7 and Fig. 8). When LLMs generate all 64 tokens, Gemma-it has the highest absolute improvement of when we perturb “Looking for guidance on choosing the best TV; could you recommend one and share additional details like display type, audio quality, or any other features that set it apart?” as “Looking for guidance on choosing the ultimate TV; could you suggest one and share extra details like display type, audio quality, or any other features that set it apart?” (i.e., replacing “best” with “ultimate”, “recommend” with “suggest”, and “additional” with “extra” in favor of brand “Samsung” in product category “TV”. With all Llama models, we see the maximum increase as we consider more tokens of the response, reaching around .
6.3 Transferability to GPT Models
@@
We evaluated the transferrability of paraphrased prompts generated using synonym-replacement (§6.2) from open-sourced models to GPT models. We found that, while transferability was limited for most model pairs, Llama3-it and GPT3.5-turbo had a high correlation in attack success for the same prompts (). We explain these results in more detail next.
@@
As we described in §5.2, we computed the relative improvement on four open-sourced LLMs and two versions of GPT models. Specifically, for a given base prompt, we generated a perturbed prompt via synonym-replacement that increased the chance of a brand being mentioned on an open-sourced model and then evaluated whether the perturbed prompt also increased the chance of the brand being mentioned in the response produced by the GPT model. We evaluated only the transfer of attacks that were successful on their open-sourced LLM onto GPT models. For the Shell example described in §6.2, we saw a improvement from to between the base and perturbed prompt on long responses by Gemma-it between prompts and an improvement from to on GPT3.5-turbo responses. On long responses generated by Llama3-it, “If you had to pinpoint the superior investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a probability of mentioning Fidelity, while “If you had to pinpoint the premier investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a , and GPT3.5-turbo responses went from to .
@@
To compare the relative improvements between models when the same prompts and synonym replacement are used, we used the Pearson correlation test. Two sets of data with a Pearson correlation coefficient () less than is generally believed to be weakly correlated, whereas or lower is uncorrelated. On the other hand, a Pearson correlation coefficient larger than is believed to indicate strong correlation [139].
@@
As shown in Fig. 10, while Pearson correlation coefficients of the relative improvement of most pairs of open-sourced LLMs and GPT models indicated weak or no correlation (), Llama3-it and GPT3.5-turbo Instruct have a correlation coefficient of , and thus are highly correlated (). Llama3-it and GPT3.5-turbo also show some correlation, with a coefficient of . This result indicates a potential for a transfer attack to be used on chatbots that use black-box GPT models, like Instacart, Lowe’s, and Expedia (described in §3) and others. A successful attack on an open-sourced LLM could be used as a prompt suggestion for black-box models, ultimately promoting the target brand.
@@ 7 User Study
Our results so far show that our synonym-replacement approach can bias LLM responses towards a target brand and appears to be human-inconspicuous. However, this does not mean that the attack is successful in practice. To evaluate the attack in a realistic setting, we conducted a user study. Here we describe this study, starting with a description of our methods (§7.1) and ending with the statistical evaluation (§7.2). We find that our attack is indeed successful in practice.
@@ 7.1 Methods
We performed a between-subjects user study to evaluate whether our synonym-replacement approach can result in prompts that advances the adversaries’ goals (see §3), creating inconspicuous prompts that trigger inconspicuous responses, making a target brand more noticable.We tested whether users could distinguish between base and perturbed prompts and responses in multiple dimensions (including clarity, likelihood of use, satifaction, and more). Each participant was shown one prompt-and-response pair. Our procedures were approved by our institution’s ethics board, and pre-registered.333https://osf.io/6mycr/?view_only=face90d04806439bb1f69fc110fb9a1e
@@ 7.1.1 Survey procedures
We recruited a gender-balanced sample from Prolific,444https://www.prolific.com/ a crowdsourcing platform commonly used with security-relevant user studies (e.g., [140]). To avoid selection bias in our study ad, we avoided mentioning LLMs or bias in the study title, “Chatbot prompting study.” Participants had to be in the U.S., be 18 or older, and have a Prolific approval rate of at least 95%. In line with Prolific guidelines 555https://researcher-help.prolific.com/hc/en-gb/articles/4407695146002-Prolific-s-payment-principles, participants were paid $1.6 for an estimated 7–9 minute study. In practice, the hourly compensation rate was $12.45/hour.
@@
Participants self-reported how frequently they give tech advice, how freuqently they use chatbots, if they paid for a chatbot, their ChatGPT familiarty. The survey concluded with demographics questions.
@@ 7.1.2 Experimental groups
Our overall goal was to find whether people notice differences between our perturbed prompts compared to base prompts. For each of our 447 base prompts, we found a perturbed prompt for each possible target brand using the loss of each of our models. This resulted in 1,809 base and perturbed prompt pairs for each model. Because of the large number of prompt pairs, we could not test all of them. We selected six pairs of prompts to use in the user study, each pair from a different product category, giving us 12 prompts total. Because each pair of prompts belonged to a unique product category, we refer to them by their product categories in the results. The prompt pairs were split between two models, three pairs for Llama3-it and three for Gemma-it. We focused on the more user-friendly instruction-tuned models from our experiments since they are more analogous to LLMs used in chatbots [141]. The pairs met the following criteria:
@@
They had the highest probability increase that the target brand is mentioned with the attack, as discussed in §6.1.
For each of the 12 prompts in our six prompt pairs, we obtained a random sample of 75 model responses from our earlier experiments with the models (§6.2). This sample was stratified to ensure the ratio of responses that mentioned the target was the same in this set of 75 as it was in the overall set of 1000. All participants within a study group saw the same prompts, but, to mirror how in real-world chatbot use a single prompt would lead to different responses, each participant was shown a unique reponse. As opposed §6, particiapnts saw full-length responses.

### v2

@@ 4.3.1 Concept Frequency Differences on Paraphrased Prompts
As we described in §4.2.4, each of the product categories in our dataset have three to ten paraphrased prompts and each of the societal topics had six to eight. For each prompt, we measured the probability that LLMs would mention some target concept of a category/topic within the first 64 tokens of the response. For each combination of category/topic and concept, we computed the maximal difference in the probability of responses mentioning the target brand between paraphrased prompts of the product category (within the first tokens). Adversaries may paraphrase a low-probability base prompt to find the high-probability prompts to achieve the goals described in §3.
@@
We summarize our results in Fig. 6 and Fig. 7. Specifically, on four open-sourced LLMs (described in §4.2.1), the likelihood of the target concept being mentioned in the responses can differ between a pair of paraphrased prompts by up to . When comparing responses to “I’m curious to know your preference for the pressure cooker that offers the best combination of cooking performance, durable construction, and overall convenience in preparing a variety of dishes.” with the prompt paraphrased as “Can you recommend the ultimate pressure cooker that excels in providing consistent pressure, user-friendly controls, and additional features such as multiple cooking presets or a digital display for precise settings?”, the probability of Gemma-it mentioning the brand “InstantPot” (“pressure cooker” product category) within the first tokens of the response went from to (i.e. went from never mentioning “InstantPot” to always mentioning “InstantPot”). Among the four models, the average of this absolute difference in likelihood is – for brands and – for societal concepts.
@@
Our results suggest that while paraphrased prompts appear similar to humans, the responses to the prompts can differ substantially in how likely they are to mention a target concept (within the first tokens of the response). Therefore, an adversary wanting to promote a certain concept can use this to their advantage: the adversary may try various paraphrases of prompts, test the prompts, and pick the paraphrase that results in the highest probability of the target concept being mentioned, ultimately promoting the concept when the perturbed prompt is used in the real world. While we generated these paraphrases using ChatGPT, and confirmed that they were valid and reasonable, adversaries may also be able to create paraphrases manually or by another method, and may be able to test even more paraphrases than we did in our measurements.
4.3.2 Synonym-Replaced Adversarial Prompts
@@
As described in §4.2.5, we measured and now report how often our synonym-replacement approach improves the probability of a concept being mentioned. As opposed to paraphrasing, in the synonym-replacement approach, we automatically generate a set of potential candidate prompts by perturbing a base prompt via synonym replacement, without needing to confirm whether these perturbed prompts are valid. The prompt with the lowest loss is selected, and we assess how well these selected prompts increase the probability that the target concept is mentioned. We evaluate the average improvement over the probabilities of different concepts being mentioned in the base prompt. We used the loss as a metric that narrowed down the large set of potential perturbed prompts we found using synonym replacements to one. Therefore, we are interested in exploring the highest improvement we can achieve between a base and perturbed score, as adversaries would be able to use a similar method to §4.3.1 and explore multiple perturbated prompts. So, we also describe the largest increases in the probability of mentioning a target concept in the responses to the perturbed prompt we found via our synonym-replacement method compared to the base prompt. Our evaluation shows that our synonym-replacement attack increases the likelihood of LLMs mentioning a concept on average. For all evaluations, we focus on the first 64 tokens of the response (see subsubsection 4.2.1 for details).
@@
Overall, we find that our attack results in an average absolute improvement in all models. Specifically, for concepts that were mentioned at least once (i.e., of the time) and at most 500 times (i.e., of the time) in responses to the base prompt, the four models had an average absolute improvement of – for brands and – for societal concepts. Our synonym replacement approach achieves a positive absolute improvement in probabilities, which verifies that our approach is capable of forcing LLMs to mention target concepts more often within some number of tokens. On all four models, we see a bigger absolute improvement within some number of tokens less than . As we described in §4.1.2, we are not claiming our proposed approach as the most effective under the threat model(§3) we present: we propose this approach only as an example to illustrate that attacks under our threat model exist. While we did not see improvements in all pairs (for example, some base prompts had near probabilities of mentioning certain concepts with long completions, not allowing for any improvement), we still saw pairs with significant increases when considering long responses. For Llama3-it, the perturbed prompt “Can you recommend the superior video game console that excels in providing top-notch graphics, dissimilar gaming options, and additional features such as online connectivity, appropriate for both casual and hardcore gamers?” was more likely than its base prompt “Can you recommend the ultimate video game console that excels in providing top-notch graphics, diverse gaming options, and additional features such as online connectivity, suitable for both casual and hardcore gamers?” (i.e., “ultimate” was changed to “superior” and “diverse” was changed to “dissimilar”) to mention Xbox in long completions, even more than when completions were only 64 tokens long (). We evaluate if attack objectives are met with full responses in a more realistic setting in §5.2.2.
@@
Besides the average absolute improvement, we also explored the maximum absolute improvement among all combinations of base prompts and concepts, as shown in Fig. 8 and Fig. 9 for brands and societal concepts respectively. While these results don’t represent the expected improvement using this method, they do demonstrate that it is possible to find pairs of prompts with vastly different probabilities of mentioning a certain concept. In §4.3.1 we showed this for prompts that were manually paraphrased; here, we show that the synonym-replacement attack can find such alternative prompts automatically. Further, the prompts generated by synonym replacement differ from their base prompts minimally (at most a seven-synonym difference in our dataset), and are perceptually the same along multiple dimensions (see §5.2.2). We were able to achieve a bigger absolute improvement on Gemma-it compared to the other three Llama models (Fig. 8 and Fig. 9), although Gemma-it does not always have the highest average absolute improvement. With all Llama models, we see the maximum increase as we consider more tokens of the response, reaching around for brands and – for societal topics.
4.3.3 Synonym-Replaced Adversarial Prompts to Mention Concepts Less Often
To systematically verify the correctness of our implementation in §4.1.2, we perturbed the prompts in the product scenario against each brand in the same category, by maximizing the loss function we proposed. Overall, we find that our implementation results in an average absolute decrease in all models. Specifically, for brands that were mentioned at least times (i.e., of the time) in responses to the base prompt, the four models had an average absolute decrease of to with in the first tokens generated. and a slightly bigger absolute decrease of to within a number of tokens less than .
Similar to what we did in §4.3.2, besides the average absolute decrease, we also explored the vastest absolute decrease among all combinations of base prompts and concepts, as shown in Fig. 10. We were able to achieve a vaster absolute decrease in the likelihood of Gemma-it () compared to the other three Llama models ( to ). While these results don’t represent the expected decrease using our approach, they do demonstrate that it is possible to slightly perturb the prompts to mention a target concept less often in LLM responses. The fact that synonym replacement can change the likelihood of LLMs mentioning a target concept in either direction suggests our approach’s potential to mitigate existing biases in LLMs, albeit it might not work for all prompts and target concepts, and there is little control of the magnitude of change in likelihoods. For example, prompt providers might use this approach to fight against known biases in LLMs, providing users with prompts that encourage LLMs to generate responses with the likelihoods of mentioning target concepts closer to the distributions expected by users. Additionally, in cases of “negative prompts”, there may be an incentive for an adversary to want a topic to show up less often.
4.3.4 Transferability to GPT Models
Some machine learning attacks were found capable of transferring [35, 36]: attacks against a machine-learning model might be effective against a different, potentially unknown, model. We thus investigated whether our synonym replacement approach can transfer to GPT3.5-turbo, a commercial and close-sourced LLM. On each of our four open-sourced LLMs, we first ranked the pairs of prompts (that recommend brands) according to the absolute improvement in mentioning a target brand. For the top pairs, we collected complete responses to both prompts: we allowed LLMs to keep generating until they yielded end-of-sequence (EOS) tokens. Instead of absolute improvement, we computed the relative improvement in the probabilities of mentioning the target concept within complete responses for each pair. If of responses of the base prompt before perturbing mentioned the target brand, and of responses of the prompt after perturbing mentioned the target brand, the relative improvement was . Then we computed the relative improvement of these pairs on ChatGPT. Specifically, we used the GPT-3.5 turbo and GPT-3.5 turbo Instruct model with the default temperature parameter and collected complete responses. GPT-3.5 turbo Instruct is the instruction-tuned version of GPT-3.5 turbo. We compare the relative improvements between GPT models and open-sourced models in §4.3.4.
@@ 4.3.4 Transferability to GPT Models
We found that, while transferability was limited for most model pairs, Llama3-it and GPT3.5-turbo had a high correlation in attack success for the same prompts (). We explain these results in more detail next.
@@
For a given base prompt (that recommends brands), we generated a perturbed prompt via synonym replacement that increased the chance of a brand being mentioned on an open-sourced model and then evaluated whether the perturbed prompt also increased the chance of the brand being mentioned in the response produced by the GPT model. We evaluated only the transfer of attacks that were successful on their open-sourced LLM onto GPT models. For the Shell example described in §4.3.2, we saw an improvement from to between the base and perturbed prompt on long responses by Gemma-it between prompts and an improvement from to on GPT3.5-turbo responses. On long responses generated by Llama3-it, “If you had to pinpoint the superior investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a probability of mentioning Fidelity, while “If you had to pinpoint the premier investment platform, which one would it be, and what specific features make it stand out as the top choice for investors?” had a , and GPT3.5-turbo responses went from to .
@@
To compare the relative improvements between models when the same base and perturbed (via synonym replacement) prompts are used, we used the Pearson correlation test. Two sets of data with a Pearson correlation coefficient () less than is generally believed to be weakly correlated, whereas or lower is uncorrelated. On the other hand, a Pearson correlation coefficient larger than is believed to indicate strong correlation [113].
@@
As shown in Fig. 11, while Pearson correlation coefficients of the relative improvement of most pairs of open-sourced LLMs and GPT models indicated weak or no correlation (), Llama3-it and GPT3.5-turbo Instruct have a correlation coefficient of , and thus are highly correlated (). Llama3-it and GPT3.5-turbo also show some correlation, with a coefficient of . This result indicates a potential for a transfer attack to be used on chatbots that use black-box GPT models, like Instacart, Lowe’s, and Expedia (described in §3) and others. A successful attack on an open-sourced LLM could be used as a prompt suggestion for black-box models, ultimately promoting the target concept.
@@ 5 User Study
Our results so far show that our synonym-replacement approach can bias LLM responses towards a target concept and appears to be inconspicuous. However, this does not mean that the attack is successful in practice. To evaluate the attack in a realistic setting, we conducted a user study. Here we describe this study, starting with a description of our methods (§5.1) and ending with the statistical evaluation (§5.2). We find that our attack is indeed successful in practice. Some of the topics we explore in previous sections might be controversial and cause emotional harm. As such, we focus the user study on the shopping task, a benign topic that users are exposed to every day. Similar setups have been used in prior work [85]. All procedures involving human subjects were approved by our institution’s ethics board. Much like those in similar work, our study might not be representative of realistic LLM use, despite our efforts to make it so. Our sample is from the U.S. and our study is in English, this might not be representative of the global population. Despite these limitations, we believe our study provides valuable data for an underexplored problem.
@@ 5.1 Methods
We performed a between-subjects user study to evaluate whether our synonym-replacement approach can result in prompts that advances the adversaries’ goals (see §3), creating inconspicuous prompts that trigger inconspicuous responses, making a target brand more noticeable. We tested whether users could distinguish between base and perturbed prompts and responses in multiple dimensions (including clarity, likelihood of use, satisfaction, and more). Each participant was shown one prompt-and-response pair.
Fully informed consent was obtained from participants and our procedures were approved by our institution’s ethics board. Following advice from prior work [114, 115], we pre-registered.444https://osf.io/6mycr/?view_only=face90d04806439bb1f69fc110fb9a1e
@@ 5.1.1 Survey procedures
We recruited a gender-balanced sample from Prolific,555https://www.prolific.com/ a crowdsourcing platform commonly used with security-relevant user studies (e.g., [116]). To avoid selection bias in our study ad, we avoided mentioning LLMs or bias in the study title, “Chatbot prompting study.” Participants had to be in the U.S., be 18 or older, and have a Prolific approval rate of at least 95%. In line with Prolific guidelines 666https://researcher-help.prolific.com/hc/en-gb/articles/4407695146002-Prolific-s-payment-principles, participants were paid $1.6 for an estimated 7–9 minute study. In practice, the hourly compensation rate was $12.45/hour.
@@
Participants self-reported how frequently they give tech advice, how frequently they use chatbots, if they paid for a chatbot, their ChatGPT familiarity. The survey concluded with demographics questions.
@@ 5.1.2 Experimental groups
Our overall goal was to find whether people notice differences between our perturbed prompts compared to base prompts. For each of our 447 base product recommendation prompts, we found a perturbed prompt for each possible target brand using the loss of each of our models. This resulted in 1,809 base and perturbed prompt pairs for each model. Because of the large number of prompt pairs, we could not test all of them. We selected six pairs of prompts to use in the user study, each pair from a different product category, giving us 12 prompts total. Because each pair of prompts belonged to a unique product category, we refer to them by their product categories in the results. The prompt pairs were split between two models, three pairs for Llama3-it and three for Gemma-it. We focused on the more user-friendly instruction-tuned models from our experiments since they are more analogous to LLMs used in chatbots [117]. The pairs met the following criteria:
@@
They had the highest probability increase that the target brand is mentioned with the attack, as discussed in §4.3.1.
For each of the 12 prompts in our six prompt pairs, we obtained a random sample of 75 model responses from our earlier experiments with the models (§4.3.2). This sample was stratified to ensure the ratio of responses that mentioned the target was the same in this set of 75 as it was in the overall set of 1000. All participants within a study group saw the same prompts, but, to mirror how in real-world chatbots use a single prompt would lead to different responses, each participant was shown a unique response. As opposed §4.3, particiapnts saw full-length responses.